15 Problem 2-27 (LG 2-9) Compute the present values of the following annuities first assuming that payments are made on the last day of the period and then assuming payments are made on the first day of the period: (Do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16)) points eBook Print References Present Value Present Value Interest Rate (Payment PaymentYears made on (Payment made on last day of period) first day of period) (Annual) 15% 848.09 9,668.26 22,022.93 71,412.54 9 15 25
Answer #1

Q1xQyVEnRykAAAAASUVORK5CYII=

Formula

PV(payment made on last days)

=Payment * ((1-(1+r)^(-n))/r)

=848.09*((1-(1+15%)^(-9))/15%)=$4046.73

=9668.26*((1-(1+8%)^(-15)/8%)=$82755.58

=22022.93*((1-(1+6%)^(-25))/6%)=$281527.19

=17412.54*((1-(1+33%)^(-6))/33%)=$177303

PV(payment made on first days)

=Payment +Payment * ((1-(1+r)^(-n))/r)

=848.09+848.09*((1-(1+15%)^(-9+1))/15%)=$4653.74

=9668.26+9668.26*((1-(1+8%)^(-15+1))/8%)=$89375.69

=22022.93+22022.93*((1-(1+6%)^(-25+1))/6%)=$298418.58

=71412.54+17412.54*((1-(1+33%)^(-6+1))/33%)=$232814